(1)Solve the formula for the specified variable.
F = 9/5C + 32 for C
(2)Solve the formula for the specified variable.
A = 1/2h(a + b) for a
(3)Solve the formula for the specified variable.
d = rt for r
(4)Solve the equation by factoring.
15x^2 + 26x + 8 = 0

Answer by vidya p(12)

1)Solve the formula for the specified variable.
F = 9/5C + 32 for C
Solution:
F = 9/5C + 32 for C
F - 32 = 9/5C
hence C = 5(F - 32)/ 9
(2)Solve the formula for the specified variable.
A = 1/2h(a + b) for a
Solution:
A = 1/2h(a + b) for a
2A/h = ( a + b)
2A/h - b = a
a= (2A-bh)/ h
(3)Solve the formula for the specified variable.
d = rt for r
Solution:
d = rt
hence r= d/t
(4)Solve the equation by factoring.
15x^2 + 26x + 8 = 0
Solution:
here a=15 , b= 26, c = 8
ac = product = 15 * 8
find the two factors of 120 , such that there sum is 26
120 = 20 * 6
26 = 20 + 6 ,
therefore: 15x^2 + 26x + 8 = 0
15x^2 + 20x + 6x + 8 = 0
5x( 3x + 4) + 2( 3x + 4) =0
( 3x + 4) (5x+2)=0
3x+4= 0 , or 5x + 2 =0
x = -4/3 or x = -2/5